# Community detection in heterogeneous signed networks

### v1

@@ 5.1 Simulated examples
500 1 0.1858(0.0063) 0.2355(0.0071) 0.2603(0.0229) -
2 0.1418(0.0082) 0.2802(0.0073) 0.1675(0.0089) -
3 0.0784(0.0073) 0.2951(0.0051) 0.1662(0.0080) -
1000 1 0.0845(0.0032) 0.1855(0.0074) 0.0970(0.0077) -
2 0.0465(0.0015) 0.2640(0.0044) 0.0882(0.0049) -
3 0.0226(0.0010) 0.2749(0.0044) 0.1208(0.0080) -
Table 6: Averaged relative probability estimation errors and their standard errors of SBBM (left) and JIM (right) over 50 independent replications in Example 3. The better performer in each scenario is boldfaced.
$n$ $\sigma^{2}$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$
500 1 0.4052(0.0063) 0.8484(0.0184) 1.6358(0.0486) 3.8576(0.0938)
2 0.3283(0.0089) 0.4778(0.0203) 1.2363(0.0253) 1.5960(0.0327)
3 0.2717(0.0092) 0.3165(0.0103) 1.1129(0.0168) 1.1200(0.0138)
1000 1 0.2725(0.0026) 0.5233(0.0055) 1.6849(0.0516) 3.7927(0.0585)
2 0.2084(0.0021) 0.2893(0.0035) 1.2639(0.0263) 1.5614(0.0211)
3 0.1751(0.0021) 0.2085(0.0024) 1.1328(0.0175) 1.1076(0.0092)
It is evident that SBBM delivers the best performance in terms of both probability estimation and community detection in almost all examples, except that its probability estimation accuracy is comparable to JIM in Example 1 with strong-balanced signed networks. Note that JIM is not applicable to detect communities in weak-balanced signed networks with $K\geq 2$, and thus the blanks in Tables 3 and 5. Its probability estimation accuracy also becomes substantially less competitive than SBBM in Examples 2 and 3 with weak-balanced signed networks. Furthermore, across all examples, the community detection errors and probability estimation errors of SBBM decrease significantly with $n$, confirming its asymptotic consistencies in Theorems 2 and 3. In contrast, the performance of SLP and SPONGE improves slowly with $n$ in Examples 2 and 3 with node heterogeneity, which is possibly due to the fact that node heterogeneity is largely ignored by SLP and SPONGE. Particularly, in Example 3, the performance of SBBM becomes better as $\sigma$ increases or node heterogeneity becomes stronger, yet such a positive trend is not observed in the performances of SLP and SPONGE. This again confirms the capability of SBBM in better capturing node heterogeneity than the other methods.
@@ 5.2 International relation network
We now turn to apply SBBM to analyze an international relation network constructed from two sources, the BACI trade database (CEPII, version 2025-01) [CEPII:2010-23], which harmonizes annual bilateral trade flows, and the Global Sanctions Data Base R4 [felbermayr2020global; yalcin2025global] that catalogs economic sanctions between economies. For each pair of economies, if economic sanctions have ever existed between them during 2015-2023, a negative edge is assigned; otherwise, a positive edge is assigned if their bilateral trade volume ranks in the top $1/30$. This pre-processing leads to a signed network with 164 economies, 580 positive edges, and 589 negative edges. It contains 2,756 strong-balanced triads, 3,559 weak-balanced triads, and 847 unbalanced triads. It is evident that weak-balanced triads occur more than four times as frequently as unbalanced ones, in sharp contrast to the expected ratio 5:3 under random assignment. This substantial deviation from randomness indicates a pronounced weak balance structure in the international relation network.
The community structure in the international relation network is then estimated via SBBM, SLP, SPONGE and JIM, respectively. Also, following the suggestion in zhang2022signed, the number of communities $K=4$ is determined by optimizing the Bayesian information criterion [saldana2017many]. For comparison, we compute the signed modularity gomez2009analysis of estimated communities from different methods. Particularly, for each $i\in[n]$, let $k^{+}_{i}=\sum_{j\in[n]}a^{+}_{ij}$, $k^{-}_{i}=\sum_{j\in[n]}a^{-}_{ij}$, $m^{+}=\frac{1}{2}\sum_{i,j\in[n]}a^{+}_{ij}$, and $m^{-}=\frac{1}{2}\sum_{i,j\in[n]}a^{-}_{ij}$. The signed modularity of a given community assignment $\{\psi_{i}\}_{i\in[n]}$ is defined as
$Q_{\text{signed}}=\frac{1}{2m^{+}}\sum_{i,j\in[n]}\left(a^{+}_{ij}-\frac{k_{i}^{+}k_{j}^{+}}{2m^{+}}\right)\delta(\psi_{i},\psi_{j})-\frac{1}{2m^{-}}\sum_{i,j\in[n]}\left(a^{-}_{ij}-\frac{k_{i}^{-}k_{j}^{-}}{2m^{-}}\right)\delta(\psi_{i},\psi_{j}),$
where $\delta(\psi_{i},\psi_{j})=1$ if $\psi_{i}=\psi_{j}$ and 0 otherwise. Note that $Q_{\text{signed}}$ takes values in [2,2][-2,2], and a larger value indicates more positive edges within communities and more negative edges between communities. Table 7 reports the signed modularities of all methods, showing that SBBM achieves the largest value, thereby demonstrating its advantages against the other methods.
@@
$Q_{\text{signed}}$ 0.4266 0.1943 0.1296 -0.0032
Figure 2 further displays the heatmap of the international relation network, where nodes are reordered according to $\widehat{\bm{Z}}$ from SBBM, and positive and negative edges are dyed into red or blue. It is clear that most positive edges occur within the same community, while negative edges occur across different communities, and a particularly high concentration of positive edges can be observed in the largest community.
@@
We also plot the estimated communities of SBBM on a world map in Figure 3. The sizes of the four estimated communities are 106, 29, 24, and 5, respectively. The largest one is centered on China and mainly consists of developing economies in Asia, Africa and Latin America. The second community is centered on the United States and mainly consists of developed economies in Europe and North America. The third community mainly consists of Afghanistan, Belarus, Iran, Syria, North Korea, Russia, Sudan and Lebanon, and the fourth community consists of Bosnia and Herzegovina, Cyprus, Eritrea, Moldova, and Ukraine. The economies in these two communities are conflict-affected or frequently subject to sanctions from other economies, and more interestingly, there are even stronger sanction ties between them. We further scrutinize the patterns of positive and negative edges via the estimated parameters. It is noteworthy that both China and the United States have high $\widehat{\beta}^{+}$, indicating their leading roles in forming strong within-community trading relations. On the other hand, both the United States and the United Kingdom have high $\widehat{\gamma}^{-}$, indicating their leading roles in imposing sanctions.
Figure 3: The four estimated communities by SBBM on the world map, where economics in different communities are dyed with different colors.
@@ 6 Summary
This paper proposes a novel signed block $\beta$-model to model the strong and weak balance structures in signed networks simultaneously. It introduces four sets of parameters that distinguish each node’s connection tendencies within its community and across different communities, which greatly facilitates community detection in signed networks with balance structures and node heterogeneity. We establish the identifiability conditions for the signed block $\beta$-model, and develop an efficient alternative updating algorithm for parameter estimation and community detection. Its asymptotic properties are also established in terms of both probability estimation and community detection, which are further supported by numerical experiments on both synthetic and real-world experiments. Note that statistical inference about the signed block $\beta$-model is also of interest but possibly hindered by the discrete feature of the community structure, and we leave it to future investigation.

### v2

@@ 5.1 Simulated examples
500 2 0.1396(0.0038) 0.2735(0.0058) 0.1781(0.0095) -
3 0.1036(0.0038) 0.2850(0.0057) 0.1796(0.0114) -
4 0.0776(0.0031) 0.2851(0.0057) 0.1689(0.0072) -
1000 2 0.0850(0.0053) 0.2616(0.0035) 0.1187(0.0142) -
3 0.0382(0.0039) 0.2780(0.0046) 0.1140(0.0088) -
4 0.0275(0.0012) 0.2754(0.0047) 0.1283(0.0079) -
It is evident that SBBM delivers the best performance in terms of community detection in almost all examples, except that it is comparable to JIM in Example 1 with strong-balanced signed networks. Note that JIM is not applicable to detect communities in weak-balanced signed networks with $K\geq 2$, and thus the blanks in Tables 3 and 5. Furthermore, across all examples, the community detection errors of SBBM decrease significantly with $n$, confirming its asymptotic consistencies in Theorems 3. In contrast, the performance of SLP and SPONGE improves slowly with $n$ in Examples 2 and 3 with node heterogeneity, which is possibly due to the fact that node heterogeneity is largely ignored by SLP and SPONGE. Particularly, in Example 3, the performance of SBBM becomes better as $\sigma$ increases or node heterogeneity becomes stronger, yet such a positive trend is not observed in the performances of SLP and SPONGE. This again confirms the capability of SBBM in better capturing node heterogeneity than the other methods.
@@ 5.2 International relation network
We now apply SBBM to analyze an international relation network constructed from two sources, the BACI trade database (CEPII, version 2025-01) [CEPII:2010-23], which harmonizes annual bilateral trade flows, and the Global Sanctions Data Base R4 [felbermayr2020global; yalcin2025global] that catalogs economic sanctions between economies. For each pair of economies, if economic sanctions have ever existed between them during 2022–2024, a negative edge is assigned; otherwise, a positive edge is assigned if their bilateral trade volume ranks in the top 5 for either economy. This pre-processing leads to a signed network with 230 economies, 947 positive edges, and 1,476 negative edges. It contains 6,451 strong-balanced triads, 10209 weak-balanced triads, and 605 unbalanced triads. It is evident that weak-balanced triads occur more than sixteen times as frequently as unbalanced ones, in sharp contrast to the expected ratio 5:3 under random assignment. This substantial deviation from randomness indicates a pronounced weak balance structure in the international relation network.
The community structure in the international relation network is then estimated via SBBM, SLP, and SPONGE with $K=3$, which is determined by the Gap statistic [tibshirani2001estimating; zhao2011community], and via JIM with $K=2$. For comparison, we compute the signed modularity [gomez2009analysis] of estimated communities from different methods. Particularly, for each $i\in[n]$, let $k^{+}_{i}=\sum_{j\in[n]}a^{+}_{ij}$, $k^{-}_{i}=\sum_{j\in[n]}a^{-}_{ij}$, $m^{+}=\frac{1}{2}\sum_{i,j\in[n]}a^{+}_{ij}$, and $m^{-}=\frac{1}{2}\sum_{i,j\in[n]}a^{-}_{ij}$. The signed modularity of a given community assignment $\{\sigma(i)\}_{i\in[n]}$ is defined as
$Q_{\text{signed}}=\frac{1}{2m^{+}}\sum_{i,j\in[n]}\left(a^{+}_{ij}-\frac{k_{i}^{+}k_{j}^{+}}{2m^{+}}\right)\delta(\sigma(i),\sigma(j))-\frac{1}{2m^{-}}\sum_{i,j\in[n]}\left(a^{-}_{ij}-\frac{k_{i}^{-}k_{j}^{-}}{2m^{-}}\right)\delta(\sigma(i),\sigma(j)),$
where $\delta(\sigma(i),\sigma(j))=1$ if $\sigma(i)=\sigma(j)$ and 0 otherwise. Note that $Q_{\text{signed}}$ takes values in [2,2][-2,2], and a larger value indicates more positive edges within communities and more negative edges between communities. Table 4 reports the signed modularities of all methods, showing that SBBM achieves the largest value, thereby demonstrating its advantages against the other methods.
@@
$Q_{\text{signed}}$ 0.7012 0.5099 0.4163 3.1717e-5
Figure 2 further displays the heatmap of the international relation network, where nodes are reordered according to $\widehat{\sigma}$ from SBBM, and positive and negative edges are dyed into red or blue. It is clear that most positive edges occur within the same community, while negative edges occur across different communities, and a particularly high concentration of negative edges can be observed between the first and second communities.
@@
We also plot the estimated communities of SBBM on a world map in Figure 3. The sizes of the three estimated communities are 90, 87, 53, respectively. The first community is centered on the United States and the European Union, encompassing most developed economies and their regional partners in the Americas. The second community is anchored by China and Russia, aggregating major emerging economies such as India, Saudi Arabia that typically do not align with Western-led sanctions. The third community forms a distinct cluster in the Asia-Pacific region, comprising Japan, South Korea, Australia, and members of the Association of Southeast Asian Nations. This tripartite structure aligns with recent findings on geoeconomic fragmentation [gopinath2025changing], confirming the divergence into U.S.- and China-centric spheres. Crucially, the result identifies an intermediate Pacific economic community that maintains regional integration despite divergent security alliances.
Figure 3: The three estimated communities by SBBM on the world map, where economics in different communities are dyed with different colors.
@@ 6 Summary
This paper proposes a novel signed block $\beta$-model to model the strong and weak balance structures in signed networks simultaneously. It introduces four sets of parameters that distinguish each node’s connection tendencies within its community and across different communities, which greatly facilitates community detection in signed networks with balance structures and node heterogeneity. We establish the identifiability conditions for the signed block $\beta$-model, and develop a two-step algorithm for parameter estimation and community detection. Its asymptotic properties are also established in terms of both parameter estimation and community detection, which are further supported by numerical experiments on both synthetic and real-world experiments. Note that statistical inference about the signed block $\beta$-model is also of interest but possibly hindered by the discrete feature of the community structure, and we leave it to future investigation.